# Pretraining Strategy for Neural Potentials

### v1

@@ 3.2 Experiment Results
RPBE EGNN 220.412 1597.309
EGNN 89.798 906.139 -58.87% -42.19% 0.25
GNS 214.970 1842.212
GNS 103.403 2217.729 -51.80% 20.31% 0.5
@@
ForceNet 231.297 - - - 0.5
Tip3p EGNN 61.481 685.071
EGNN 39.291 241.853 -36.09% -64.70% 0.9
GNS 52.021 638.216
GNS 49.104 540.406 -5.61% -15.33% 0.5
@@
Table 1: Quantitative force and energy performance analysis on RPBE and Tip3p datasets. Performance improvements are calculated between pretrained models and their trained-from-scratch counterparts. The units for force and energy RMSE are ($meV/$\mathrm{\SIUnitSymbolAngstrom}$$) and ($meV$).
Following the described training setup, we conducted experiments to investigate if masking pretraining would help the finetuning objective over energy and force prediction accuracy on molecule potential datasets. All the configurations are trained on four different seeds, and we report the averaged results. As shown in Table 1, we compared the RMSE of predictions of force and energy for our baseline models. The results show that our proposed masking pretraining strategy provides a significant improvement over both objectives across datasets. Most of the RMSE objectives are reduced by pretraining with our proposed masking pretext task. The average improvement of force and energy RMSE for EGNN and GNS is $38.09\%$ and $25.48\%$. Particularly, the pretrained EGNN model 63 gained more than $35\%$ performance improvement on all objectives. Masking pretraining also helps the GNS model 64 reduce $28.71\%$ force RMSE. This implies that our approach is model-agnostic (equivariant and covariant models) and can be generalized to both force-centric models and energy-centric models.
It is worth noting that the ForceNet model65 is hardly converging on both datasets without pretraining. Since ForceNet’s results are several orders of magnitude higher than other models’ results, we did not report out-of-scale results for clarity. However, we observe that finetuning ForceNet becomes more stable after pretraining in comparison to ForceNet trained from scratch. We also observe a similar trend that masked pretraining can effectively stabilize the training of other models. We argue that masked pretraining could facilitate more stable training and faster convergence. We elaborate on this point with training logs in the supplemental information section 6.2.
@@ 3.3 Comparison over models trained from scratch with additional epochs
Extended experiments are also conducted to help demonstrate the efficacy and necessity of our two-stage training approach in comparison to simply extending the training duration without pretraining. In contrast to trained from scratch, pretrained models have yielded significant improvements on the RPBE dataset, as shown in figure 2. In the first row of figure 2, We compared the energy and force RMSE of our pretrained models to the models trained from scratch for either regular 300 or an extended 400 epochs. The results show that our training strategy of 100 epochs of pretraining followed by 300 epochs of finetuning still outperforms the elongated train-from-scratch models.
The performance gap in force RMSE between EGNN and GNS models trained on the RPBE dataset with and without pretraining sustains when compared to an extended training duration. It underscores the effectiveness of pretraining over merely increasing the training iterations. On Tip3p, the difference between the pretrained models and the model trained from scratch with 75 epochs is reduced on force RMSE. Energy-wise, a similar trend is observed. It is worth noting that the pretrained EGNN on the RPBE dataset achieves a higher energy RMSE improvement when compared with EGNN trained from scratch with 400 epochs instead of 300 epochs. We attribute it to the overfitting of the model to this dataset. This observation indicates that for models converging on smaller datasets given certain epochs, starting from a point where the model has learned meaningful molecule featurization (with pretraining), the network is less likely to converge to worse local minima during the final training phase.
Pretraining gives the models a better initialization and facilitates models to extract meaningful representation from the unlabelled data, resulting in more adept and competent models. Therefore, our findings suggests the integration of pretraining as an essential component in building machine learning potential training pipelines.
@@ 3.4 Comparison with Denoising
Figure 3: Force and Energy performance comparison between models pretrained with masking and denoising tasks
@@
We provide the comparative performance metrics of our proposed masking and denoising methods, which is critical in evaluating the efficacy of computational models in predicting molecular properties. Figure 3 provided a direct comparison between these two methods across two datasets, for force and energy predictions. To create fair comparisons, we held masking pretrained models’ results from table 1 constant. We implemented denoising as described in Sec 2.3. The mean and standard deviation of the Gaussian Noise added are set to be 0 and 0.5 throughout the experiments.
@@
We observe that in the RPBE dataset, the performance of denoising is similar and competitive to that of our method. We attribute this observation to the effectiveness of the denoising approach in single molecule and attribute label pairs datasets such as ISO17, and ANI-1x as demonstrated in recent research 57, 60. Containing on average 27 water molecules, data points in the RPBE dataset are comparatively smaller than those in the Tip3p dataset. Importantly, denoising has shown difficulties scaling to the Tip3p dataset. EGNN model under the masking condition has a energy RMSE of about 241.85 $meV$. The performance deteriorates significantly when denoising is applied, with poor convergence and RMSE rising to approximately 2442.75 $meV/$\mathrm{\SIUnitSymbolAngstrom}$$ by the end of the finetuning stage. Denoising also exhibits degraded performance on the Energy RMSE metric when pretraining using the GNS model on the Tip3p dataset. Both EGNN and GNS pretrained using masked pretraining have shown stable convergence during training without suffering from scalability and achieved superior performance.
@@ 4 Conclusion
We present a new method of pretraining for improving the performance of GNNs on forcefield learning. We demonstrate that our method could achieve better accuracy on both force and energy prediction. Moreover, it also facilitates more stable training and exhibits better convergence in comparison to models without pretraining or pretrained with other pretext tasks such as denoising. These results indicate the effectiveness of our method for further utilization of GNNs to facilitate the development of neural potentials. Scaling our method with larger and more advanced GNNs, training on larger-scale datasets and molecular systems, and extending our methods to auto-regressive models are promising directions for future work.

### v2

@@ 3.2 Experiment Results
RPBE EGNN 221.0 (81.3) 1570.1(660.7)
EGNN 90.9(8.3) 907.6(96.1) -58.9% -42.2% 0.25
GNS 214.2(74.8) 1842.2(616.4)
GNS 103.2(1.3) 2216.4(76.9) -51.8% 20.3% 0.5
@@
ForceNet 231.3(5.1) - - - 0.5
Tip3p EGNN 61.5(4.0) 685.1(55.8)
EGNN 39.3(2.0) 241.9(27.3) -36.1% -64.7% 0.9
GNS 52.0(0.2) 638.2(11.6)
GNS 49.1(0.3) 540.4(8.1) -5.6% -15.3% 0.5
@@
Table 1: Quantitative analysis of force and energy RMSE on RPBE and Tip3p datasets. Changes in RMSE are shown for pretrained models compared to their trained-from-scratch counterparts, with negative percentages indicating improvements. The units for force and energy RMSE are ($meV/$\mathrm{\SIUnitSymbolAngstrom}$$) and ($meV$), respectively. All results represent the average of four runs with different random seeds, with the corresponding standard errors reported in parentheses.
Following the described training setup, we conducted experiments to investigate if masked pretraining would help the finetuning objective over energy and force prediction accuracy. All the configurations are trained four times with different random seeds, and we report the averaged results. As shown in Table 1, we compared the Root mean square error (RMSE) of predictions of force and energy. The results show that the masked pretraining strategy provides a significant improvement over both objectives across datasets. Most of the RMSE objectives are reduced by pretraining with our proposed masking pretext task. The average improvement of force and energy RMSE for EGNN and GNS is $38.09\%$ and $25.48\%$. Particularly, the pretrained EGNN model gained 47.48% and 53.45% of performance improvement on force and energy objectives respectively. Masking pretraining also helps the GNS model reduce $28.71\%$ of force RMSE. This implies that our approach is model-agnostic (equivariant and covariant models) and can be generalized to both force-centric models and energy-centric models.
It is worth noting that the ForceNet model is hardly converging on both datasets without pretraining. Since ForceNet’s results are several orders of magnitude higher than other models’ results, we did not report out-of-scale results for clarity. However, we observe that finetuning ForceNet becomes more stable after pretraining in comparison to ForceNet trained from scratch. We also observe a similar trend that masked pretraining can effectively stabilize the training of other models, and we argue that masked pretraining could facilitate more stable training and faster convergence. We further demonstrate and elaborate on this point with training loss curves of the ForceNet model in the supplemental information Section 2.
@@ 3.3 Comparison over models trained from scratch with additional epochs
Extended experiments are also conducted to help demonstrate the efficacy and necessity of our two-stage training approach in comparison to simply extending the training duration without pretraining. In contrast to trained from scratch, pretrained models have yielded significant improvements on the RPBE dataset, as shown in figure 2. In the first row of figure 2, We compared the energy and force RMSE of our pretrained models to the models trained from scratch with either regular 300 or an extended 400 epochs. The results show that our training strategy of 100 epochs of pretraining followed by 300 epochs of finetuning still outperforms the elongated train-from-scratch models.
On the RPBE dataset, the performance gap in force RMSE between EGNN and GNS models trained with and without pretraining sustains, despite the additional finetuning time. It underscores the effectiveness of pretraining over merely increasing the training iterations. On Tip3p, the difference in force RMSE between the pretrained models and the model trained from scratch with extended 75 epochs is reduced, and a similar trend is observed energy-wise. It is worth noting that the pretrained EGNN on the RPBE dataset achieves a higher energy RMSE improvement when compared with EGNN trained from scratch with 400 epochs instead of 300 epochs. We attribute it to the overfitting of the model to this dataset. This observation indicates that for models converging on smaller datasets given certain epochs, starting from a point where the model has learned meaningful molecule featurization (with pretraining), the network is less likely to converge to worse local minima during the final training phase.
Pretraining gives the models a better initialization and facilitates models to extract meaningful representation from the unlabelled data, resulting in more adept and competent models. Therefore, our findings suggest integrating masked pretraining into machine learning potential training pipelines is non-trivial and of great potential.
@@ 3.4 Comparison with Denoising
Figure 3: Force and Energy performance results from models pretrained with masking and denoising tasks on the RPBE and Tip3p datasets
@@
We provide the comparative performance metrics of our proposed masked and denoising methods, which is critical in evaluating the efficacy of different pretraining strategies in predicting molecular properties. Figure 3 provided a direct comparison between these two methods across two datasets for force and energy predictions. We held masking pretrained models’ results from table 1 constant. We implemented denoising as described in Sec 2.3. The mean and standard deviation of the Gaussian Noise added are set to be 0 and 0.5 throughout the experiments.
@@
We observe that in the RPBE dataset, the performance of denoising is similar and competitive to that of our method. We attribute this observation to the effectiveness of the denoising approach in single molecule and attribute label pairs datasets such as ISO17, and ANI-1x as demonstrated in recent research 62, 65. Containing on average 27 water molecules in each sample, the number of samples in the RPBE dataset is 77.17% less than that of the Tip3p dataset. Importantly, denoising has shown difficulties scaling to the Tip3p dataset. EGNN model with masked pretraining has an energy RMSE of about 241.85 $meV$. The performance deteriorates significantly when denoising is applied, with poor convergence and RMSE rising to 2442.75 $meV/$\mathrm{\SIUnitSymbolAngstrom}$$ by the end of the finetuning stage. Denoising also exhibits degraded performance on the Energy RMSE metric with the GNS model pretrained on the Tip3p dataset. Both EGNN and GNS pretrained using masked pretraining have shown stable convergence during training without suffering from scalability and achieved superior performance.
3.5 Experiment on Revised MD17 Dataset
We additionally evaluated our masked pretraining method with EGNN on the Revised MD17 dataset 84, which is a refined version of the MD17 dataset 85, 16. It contains molecular dynamics (MD) trajectories of 10 types of small organic molecules at DFT accuracy, each comes with 100,000 samples (except for Azobenzene which has 99,888 samples).
Following the setup in Schütt et al. 33, the training sample size is 50,000, with 1,000 samples for evaluation, and the rest for testing. For Pretraining, we combine all the training samples for each molecule as the pretraining dataset and further finetune the models on the individual molecules. Thus, the overall pretraining dataset is of size 500,000 for 10 molecules. After pretraining, the pretrained model checkpoint is subsequently finetuned on each molecule individually with the same data splits to ensure no data leakage.
We adopt the number of pretraining epochs to 15 and finetuning epochs to 200 in addition to a batch size of 32. Following prior works, we use Adam Optimizer with a learning rate of 1e-4 and a minimum learning rate of 5e-5. Prior works 86, 87 regard force and energy weights as tunable hyperparameters during training. In this work, we set them to 40 and 15 respectively, with which we observed reasonable convergence of EGNN on the Revised MD17 dataset.
We pretrained EGNN models using three masking ratios (0.25, 0.5, 0.75) and then fine-tuned them on the Uracil molecule as validation. Although all three models showed reasonable performance improvements, we transferred the model weights pretrained with a masking ratio of 0.5 and finetuned each molecule in the dataset on top of it for its best performance among all.
Molecule Pretrain Force MAE Energy MAE Change (%)
Force Energy
Revised Uracile 22.0 11.8
19.0 10.7 -13.6% -9.3%
Revised Benzene 14.6 5.4
7.7 6.2 -47.3% 14.8%
Revised Naphthalene 14.6 30.3
18.5 14.7 26.7% -51.5%
Revised Aspirin 46.7 25.2
29.5 16.0 -36.8% -36.5%
Revised Malonaldehyde 32.9 14.4
11.7 8.3 -64.4% -42.4%
Revised Ethanol 16.8 8.8
10.7 4.7 -36.3% -46.6%
Revised Toluene 17.9 13.7
12.1 10.2 -32.4% -25.5%
Revised Salicylic Acid 27.5 25.9
17.6 13.2 -36.0% -49.0%
Table 2: Force MAE and Energy MAE performance change results for EGNN with or without masked pretraining on the Revised MD17 dataset. The units for force and energy MAE are ($meV/$\mathrm{\SIUnitSymbolAngstrom}$$) and ($meV$).
We present the results for Force and Energy mean absolute error (MAE) in Table 2, showing an average reduction of 30.02% and 30.58%, respectively. These findings suggest that our method can effectively generalize to small organic molecules with diverse types of atoms that are heavier than the Hydrogen atom. A major advantage of our proposed strategy is that it can be trained in a diverse set of unlabelled configurations, after which the pretrained model can be transferred to different types of molecules. Additionally, initializing from the weights that have already learned general patterns from extensive data, the convergence of the pretrained model is also faster, allowing the model to quickly adapt to different molecular structures. Most importantly, this approach ensures consistent and reliable performance across different molecules, which is vital for robust and accurate potential energy approximation.
@@ 4 Conclusion
We present a new method of masked pretraining for improving GNN’s performance on forcefield learning. We demonstrate that our method could achieve better accuracy on both force and energy prediction. Moreover, it also facilitates more stable training and exhibits better convergence in comparison to models without pretraining or pretrained with other pretext tasks such as denoising. These results indicate the effectiveness of our method for further utilization of GNNs to facilitate the development of neural potentials. Despite our method being currently limited to molecules that contain Hydrogen atoms, we believe that exploring masking other atoms or masking a combination of different types of atoms and scaling our method with larger and more advanced GNNs, are promising directions for future work.
5 Supplementary Material
Supplementary material includes the following sections:
  • Section 1: Ablation study of the choice of masking ratios;
  • Section 2: Further analysis of the ForceNet70 results mentioned in Sec 3.2;
  • Section 3: Computational cost of masked pretraining;